Q.Choose the correct answer.
Q. The initial value value problem dy/dx=|y|1/2 ,y(0)=0 has:
a) unique solution
b) No solution
c) Infinitely many solution
d) Two solutions.
1
Expert's answer
2019-07-10T13:31:30-0400

 The initial value value problem dy/dx=|y|1/2 ,y(0)=0 has

d) Two solutions.

Find the general solution of


"y'=\sqrt{|y|}"

"\frac{dy}{dx}=\sqrt{|y|}"

1) Let "y(x) \leq 0" :


"\frac{dy}{dx}=\sqrt{-y}"

We can separate the variables:

"\frac{dy}{\sqrt{-y}}=dx"

"-\frac{d(-y)}{\sqrt{-y}}=dx"

Now, integrate the left-hand side dy  and the right-hand side dx

"-\int\frac{d(-y)}{\sqrt{-y}}=\int dx \\"

"-2\sqrt{-y}=x+C_1 \\"

"\sqrt{-y}=-\frac{1}{2}(x+C_1)"


"-y=\frac{1}{4}(x+C_1)^2, x+C_1\leq 0"

"y=-\frac{1}{4}(x+C_1)^2, x+C_1\leq 0 \text{ } (1)"


2) Let "y( x)\geq 0" :



"\frac{dy}{dx}=\sqrt{|y|} \Rightarrow \frac{dy}{dx}=\sqrt{y}"



We can separate the variables:

"\frac{dy}{\sqrt{y}}=dx"

Now, integrate the left-hand side dy  and the right-hand side dx

"\int\frac{d(y)}{\sqrt{y}}=\int dx \\"


"2\sqrt{y}=x+C_1 \\"

"\sqrt{y}=\frac{1}{2}(x+C_1)"

"y=\frac{1}{4}(x+C_1)^2, x+C_1\geq 0 \text{ } (2)"

So there is our general solution (from (1), (2)): 



"y=\frac{1}{4}(x+C_1)^2, x+C_1\geq 0\\
y=-\frac{1}{4}(x+C_1)^2, x+C_1\leq 0"

To find the particular solution, substitute the initial condition values to obtain

"y(0)=0 \Rightarrow 0=\frac{1}{4}(0+C_1)^2, \Rightarrow C_1=0"

So, the particular solution that satisfies the initial condition is

"y=-\frac{1}{4}x^2, x \geq0 \\
y=-\frac{1}{4}x^2, x \leq 0"

But the solution to equality is "y = 0" , which also satisfies the condition "y(0)=0" .


The initial value value problem dy/dx=|y|1/2 ,y(0)=0 has Two solutions:



"1)\text{ } y=\frac{1}{4}x^2, x \geq0 \\
y=-\frac{1}{4}x^2, x \leq 0"

And

"2) \text{ } y= 0 ."